OpenStudy (anonymous):

Solve the equation by factoring I ( quadratics) 3x^2+3x=18 I got x=-2, x=3 Is that correct?

OpenStudy (anonymous):

First get the 18 on the same side as the x's.

OpenStudy (anonymous):

start by subracting 18 from both sides to get 3x`2+3x -18=0

OpenStudy (anonymous):

You're almost right, but you have to watch your signs.

OpenStudy (anonymous):

I got 3x^2+3x-18=0 (3x+6)(x-3) 3x+6=0 3x=-6 x=-2 x-3=0 x=3

OpenStudy (anonymous):

That is incorrect, because you would get -3x....you want a +3x

OpenStudy (anonymous):

You can, however, factor a 3. 3(x^2 + x - 6) That factors very nicely.

OpenStudy (anonymous):

Let me try it that way

OpenStudy (anonymous):

i got the same answer

OpenStudy (anonymous):

Ok, how does x^2 + x - 6 factor?

OpenStudy (anonymous):

I think I am factoring incorrectly (x+2)(x-3)

OpenStudy (anonymous):

You are. FOIL that out and see what you get. To get a POSITIVE x, the LARGER constant must remain with the (+) sign.

OpenStudy (anonymous):

The factors are (x-2)(x+3)

OpenStudy (anonymous):

x=2 amd x=-3 ?

OpenStudy (anonymous):

Correct, THOSE are the zeroes/roots of the function.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Np